### v1

@@ 3 Methods
3.1 Sample preparation
Several clay types: bentonite, montmorillonite, illite, kaolinite, and silt, were purchased from Merck and Nagra in the form of powder and grains. The samples were preliminary milled to the sub-micron sizes. Then, clay samples were washed out through a multi-step cleaning process involving repeated dissolution, natural sedimentation preselection, washing and centrifugation cycles, dialysis, and sonication, with zero chemicals used. The resulting product was a nonprecipitating colloidal solution of clay nanoparticles in distilled water with a concentration of about 0.1 mg/ml, a bulk electrical resistivity of approximately 18 M, and a pH of 7. An electrochemically exfoliated graphene solution in distilled water was purchased from XFNano and used as received. Milli-Q water, with an 18 M resistivity, served as the water source. Polymer-filter substrates with nanopores ranging from 20 to 100 nm were purchased from Merck and used without further modification.
@@ 3.6 Scanning transmission electron microscopy
Scanning transmission electron microscopy (STEM) experiments were realized on an aberration-corrected FEI Titan Themis3 microscope operated at 300 kV. An electron probe of an 8 mrad half-convergence angle was used in association with a current of 10 pA. A dwell time of 2 µs was used for STEM imaging, which limited the total electron dose deposited on the sample for acquiring atomic-resolution STEM images to 460 e-/. Annular dark-field (ADF) and integrated differential phase contrast (iDPC) STEM images were collected simultaneously, using a collection angle ranging from 17 to 104 mrad for ADF and from 4 to 15 mrad for iDPC. iDPC STEM imaging allowed us to visualize the atomic structure of the crystal composed of light elements [50], with the moderate electron dose used to avoid degradation of the crystal structure during imaging.

### v3

@@ 4 Methods
4.1 Samples preparation and cleaning
Several clay types: bentonite, montmorillonite, illite, kaolinite, and silt, were purchased from Merck and Nagra in the form of powder and grains. The samples were carefully cleaned before use. To do this, the samples were preliminary milled to the sub-micron sizes. Then, clay samples were washed out through a multi-step cleaning process involving repeated dissolution, natural sedimentation preselection, washing and centrifugation cycles, dialysis, and sonication, with zero chemicals used. The resulting product was a nonprecipitating colloidal solution of clay nanoparticles in distilled water with a concentration of about 0.1 mg/ml, a bulk electrical resistivity of approximately 18 M, and a pH of 7. An electrochemically exfoliated graphene solution in distilled water was purchased from XFNano and used as received. Milli-Q water, with an 18 M resistivity, served as the water source. Polymer-filter substrates with nanopores ranging from 20 to 100 nm were purchased from Merck and used without modification.
@@ 4.6 Scanning transmission electron microscopy
Scanning transmission electron microscopy (STEM) experiments were realized on an aberration-corrected FEI Titan Themis3 microscope operated at 300 kV. An electron probe of an 8 mrad half-convergence angle was used in association with a current of 10 pA. A dwell time of 2 s was used for STEM imaging, which limited the total electron dose deposited on the sample for acquiring atomic-resolution STEM images to 460 e-/. Annular dark-field (ADF) and integrated differential phase contrast (iDPC) STEM images were collected simultaneously, using a collection angle ranging from 17 to 104 mrad for ADF and from 4 to 15 mrad for iDPC. iDPC STEM imaging allowed us to visualize the atomic structure of the crystal composed of light elements [39], with the moderate electron dose used to avoid degradation of the crystal structure during imaging.